OpenStudy (anonymous):

Solve the quadratic equation x2 - 4x + 57 = -5 by completing the square.

OpenStudy (anonymous):

\[x ^{2}-4x+57=-5\]

OpenStudy (anonymous):

\[x ^{2}-4x=-62\]

OpenStudy (anonymous):

\[(\frac{ -4 }{2 })^{2}=4\]

OpenStudy (anonymous):

\[x ^{2}-4x+4=-58\]

OpenStudy (anonymous):

\[\sqrt{x ^{2}}+\sqrt{4}=\sqrt{-58}\]

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

Im stuck here

OpenStudy (anonymous):

\[A.x=1\pm i \sqrt{50}\]

OpenStudy (anonymous):

\[B.x=2\pm \sqrt{58}\]

OpenStudy (anonymous):

\[C.x=2\pm i \sqrt{58}\]

OpenStudy (anonymous):

\[D.x=2\pm \sqrt{62}\]

OpenStudy (anonymous):

yeah you really are stuck

OpenStudy (anonymous):

\[x ^{2}-4x=-62\] \[(x-2)^2=-62+4=-58\] \[x-2=\pm\sqrt{52}i\] \[x=2\pm\sqrt{58}i\]

OpenStudy (anonymous):

Wow Thanks lol fan testimonial was only way of pming you Thanks alot

OpenStudy (anonymous):

yw method i hope is more clear

OpenStudy (anonymous):

It is